### v1

@@ 2 Radio emission mechanisms and causes of variability
where $I_{\nu}$ is the specific intensity (measured in ${\rm erg\,s}^{-1}\,{\rm cm}^{-1}\,{\rm Hz}^{-1}\,{\rm sr}^{-1}$ or equivalent), the frequency is $\nu$, $k_{B}$ is Boltzmann’s constant, and $c$ is the speed of light. For a source with flux density $S_{\nu}$ and solid angle $\Omega$, we have $I_{\nu}=S_{\nu}/\Omega$. Note that this form assumes radiation is in the Rayleigh-Jeans limit, with $h\nu\ll k_{B}T$, even if that is not actually true. For non-thermal sources the brightness temperature, $T_{B}$, is different from the actual electron temperature $T_{e}$, and typically $T_{B}\gg T_{e}$.
@@ 2.1.1 Accretion and jet ejection
AGN dominate the radio sky at flux densities above a few millijanskys (at centimetre wavelengths), and most show little variability (hovatta_statistical_2007; hodge_millijansky_2013; stewart_lofar_2016; mooley_caltech-nrao_2016; bell_murchison_2019; murphy_askap_2021), although fluctuations of a few to few tens of percent are not uncommon (e.g. falcke_radio_2001; mundell_radio_2009) and variations are both more significant and faster at higher radio frequencies (e.g. ackermann_radiogamma-ray_2011; hovatta_statistical_2007), especially for sub-classes of AGN such as blazars (max-moerbeck_ovro_2012; richards_blazars_2011).
@@ 2.3 Stellar emission: gyrosynchrotron, plasma, and electron cyclotron maser emission
In a number of stellar systems, especially for cooler, later-type stars we can detect gyrosynchrotron emission, which is an incoherent, (usually) non-thermal process related to synchrotron emission (dulk_radio_1985; gudel_stellar_2002). Here, however, the electrons are only mildly relativistic. This leads to broad-band emission that has significant (but not 100%) circular polarisation at higher frequencies, where it is optically thin. At lower frequencies we see a typical optically-thick spectral index of $\alpha=+2.5$ and an steep spectrum on the higher frequency side, with the power-law index depending on the underlying distribution of electron energies.
@@ 3.1.1 Radio detection of gamma-ray bursts
The first GRB for which a radio afterglow was detected was GRB 970508 (frail_radio_1997). High variability was observed in the first month after the explosion, which was attributed to diffractive scintillation, Since diffractive scintillation occurs only when the source is smaller than a characteristic size, the transition from the diffractive to refractive regime could be used to measure the angular size of the fireball, determining that it was a relativistically expanding explosion. The subsequent monitoring of this event in radio (frail_450_2000) demonstrated the utility of radio observations in measuring the calorimetric properties of GRBs.
@@ 3.1.3 Short gamma-ray bursts
Neutron star mergers have long been believed to be the origin of short gamma-ray bursts (eichler_nucleosynthesis_1989; kochanek_gravitational_1993; although it can be more complicated, as in troja_nearby_2022; rastinejad_kilonova_2022). Like with long GRBs there is a collimated, relativistic jet that impacts the surrounding medium, and so many of the phenomena related to long GRBs could also take place in in short GRBs (granot_gamma-ray_2014; berger_short-duration_2014).
@@ 3.1.4 Neutron star mergers
While a burst of gamma-rays was seen from GW170817/GRB170817A (goldstein_ordinary_2017), it was considerably lower luminosity that from most GRBs. The distinction is that the observer needs to be on-axis to see a GRB, while the more general neutron star merger phenomena could be observed off-axis, as was the case with GW170817/GRB170817A. Most subsequent searches for neutron star mergers at radio wavelengths still rely on multi-messenger triggers. Given the uncertain nature of many of those triggers and the wide sky areas they point to (e.g. petrov_data-driven_2022), even triggered searches still retain some aspects of wide-field radio surveys, with many sources to consider and tight control over false positives required. With these triggered searches, non-detections can still be used to constrain the geometry, energetics, and environments of neutron star mergers (gulati_constraints_2025). But even without such a trigger, the collimated afterglow from a SGRB/neutron star merger could still be detectable by wide-field radio surveys (metzger_extragalactic_2015; hotokezaka_radio_2016; lin_detectability_2020), although as with many similar events it may be a challenge to distinguish them from other related synchrotron transients. Nonetheless, they are promising targets to understand the diversity in jet launching and how neutron star mergers are tied to their environments. If detected, followup observations including astrometry and searches for scintillation can then be used to determine the jet geometry and energetics (dobie_constraining_2020), especially with next-generation radio facilities (dobie_radio_2021; corsi_radio_2024).
@@ 3.2 Supernovae
The simple picture of steady spherical mass-loss is, in many case, contradicted by detailed radio observations. This has been studied in detail in the case of SN 1987A in the Large Magellanic Cloud (staveley-smith_structure_1993; mccray_remnant_2016; petruk_polarized_2023), through multi-wavelength modeling of the evolving morphology and spectrum, but usually this level of detail is not possible. Instead observations can usually only constrain the flux density, and in small numbers of cases the size and rough shape (with very long baseline interferometry; e.g. demarchi_radio_2022). But even the basic lightcurve information is complex, and can show multiple radio peaks (anderson_peculiar_2017; palliyaguru_vlbi_2021; rose_late-time_2024) indicating episodic mass-loss, or broadened peaks (soderberg_radio_2005; sfaradi_dense_2024) showing aspherical ejecta. In particular, we highlight the use of unbiased surveys (stroh_luminous_2021; rose_late-time_2024) in systematically probing for unexpected emission of a large, diverse sample of SNe, and further searching for (likely) SNe that were not identified previously (gal-yam_radio_2006; dong_transient_2021).
@@
Aside from core-collapse SNe, thermonuclear SNe (Type Ia; blondin_type_2024) are of considerable interest for their elemental enrichment patterns (dwek_iron_2016) and their use in cosmology (riess_observational_1998; perlmutter_measurements_1999), but the underlying natures of their progenitors are still poorly understood (maoz_observational_2014; liu_type_2023). Radio observations can in principle distinguish between progenitor scenarios (horesh_early_2012; perez-torres_constraints_2014), constraining the presence of a non-degenerate companion. However, despite considerable investment of observational resources (e.g. panagia_search_2006; hancock_visibility_2011; chomiuk_evla_2012; margutti_inverse_2012; horesh_early_2012; perez-torres_constraints_2014) no type Ia SNe were detected for many years. This changed with the detection of of SN2020eyj by kool_radio-detected_2023, whose radio properties show signs of interaction between a blast wave and the circumstellar medium (CSM), suggesting a non-degenerate companion. However, the optical/infrared signature of this SN is also somewhat unusual, showing signs of CSM interaction (Type Ia-CSM; silverman_type_2013), so it remains to be seen whether this signature is borne out by future observations of a larger sample of Ia SNe.
@@ 3.2.1 Luminous fast blue optical transients
Sitting at the intersection of core-collapse SNe and GRBs, there are expected to be a variety of relativistic explosions that would lack high-energy emission (e.g. meszaros_tev_2001). Such sources have begun to be discovered in increasing numbers by large optical surveys. The first of these from an emerging new class is AT2018cow (prentice_cow_2018; perley_fast_2019; ho_at2018cow_2019; margutti_embedded_2019; nayana_ugmrt_2021), which was notable for having bright, fast, blue optical emission along with very bright and long-lasting (sub)millimetre emission transitioning, along with quickly declining centimetre emission whose properties are not easily reconciled with the millimetre emission (margutti_embedded_2019; ho_at2018cow_2019; ho_luminous_2022) in the usual way for synchrotron explosions. These sources are called ‘(luminous) fast blue optical transients’ or (L)FBOTs, with LFBOTs those with the highest peak luminosities ($10^{44}\,{\rm erg\,s}^{-1}$) that are achieves in the shortest times (few days), although FBOTs can span a range of lower luminosities and slower timescales (e.g. drout_rapidly_2014).
@@ 3.4 Tidal disruption events
Tidal disruption events (TDEs; hills_possible_1975; rees_tidal_1988; komossa_tidal_2015; see gezari_tidal_2021 for a recent review) occur when a star is disrupted while passing with the tidal radius of a supermassive black hole (note that other TDE scenarios involving compact objects and lower-mass black holes have been proposed, e.g. krolik_swift_2011). This leads to emission potentially across the electromagnetic spectrum, including at radio wavelengths (van_velzen_radio_2011; giannios_radio_2011; see alexander_radio_2020 for a recent review).
@@ 3.5.5 Late F, G & K type main sequence stars
Finally, some F, G & K stars exhibit coherent bursts from a range of mechanisms (e.g. davis_detection_2024) — although in general the stellar equivalents of solar radio bursts have been primarily studied on K & M dwarfs due to their extremely strong magnetic fields, since these result in stronger and more frequent flare activity (see Section 3.5.7).
@@ 3.7 X-ray binaries
Low-mass X-ray binaries, where neutron stars or black holes accrete from stellar-mass companions, are highly variable sources seen across the electromagnetic spectrum (fender_overview_2014; van_den_eijnden_new_2021; fender_jets_2006). The jets produced by these systems inject energy into the surrounding medium (gallo_dark_2005), regular star formation and ionising the surrounding gas (fender_energization_2005), and can be used to study the formation of jets and accretion across many orders of magnitude in accretor mass (from stellar-mass systems to supermassive black holes; gallo_universal_2003; gallo_radiox-ray_2014; van_den_eijnden_new_2021), giving insight into the fundamental mechanisms of jet launching (mcclintock_black_2014) and synchrotron-emitting regions (e.g. chauhan_broadband_2021). Detailed imaging of the radio jets can also show direct evidence for relativistic jet launching and superluminal motion (mirabel_superluminal_1994; fender_ultra-relativistic_2004).
@@
Observing such systems over time and across wavelengths, there is evidence for correlation between the radio and X-ray luminosities that extends across many orders of magnitude in mass (merloni_fundamental_2003; migliari_jets_2006; gallo_universal_2003)) when the X-ray emission is in the ‘hard’ state, but which may also break into multiple tracks with different levels of radio emission (soleri_nature_2011) whose origin and significance are not understood (gallo_hard_2018). However, when the X-ray emission is in other spectral states (‘intermediate’ or ‘soft’; belloni_states_2010) this correlation disappears as steady jet emission stops and discrete blobs may be ejected (fender_towards_2004). The transitions between these states likely represent disc instabilities changing the rate at which material moves from the disc onto the compact object (coriat_revisiting_2012). Such transitions can give rise to transient radio jets, either going from the ‘soft’ X-ray state to the ‘hard’ or the other way, when newly-launched jets can impact the material from previous ejection cycles (koljonen_2006_2013).
@@ 3.9 Long period transients and Galactic Centre Radio Transients
LPTs also share some radio properties with radio pulsating white dwarf binary systems (that consist of a white dwarf star with an M-dwarf companion) [^8] Although these are sometimes referred to in the literature as ‘white dwarf pulsars’, we find this terminology confusing as it (a) it mixed both physical and phenomenological information, and (b) ‘pulsar’ is a well established term for neutron stars; which are not present in these systems.. There have been potentially six such systems detected to date: AR Sco, with a radio period of $\sim 1.95$ min in a 3.5 h orbit (marsh_radio-pulsing_2016); J1912-4410, with a radio period of $\sim 5.3$ min in a 4 h orbit (pelisoli_53-min-period_2023); ILT J1101++5521, with a much longer period of 2.1 h, matching the orbital period (de_ruiter_transient_2024); GLEAM-X J0704-37, with a period of 2.9 h (again matching the orbit; hurley-walker_29_2024; rodriguez_spectroscopic_2025); SDSS J230641.47+244055.8 with pulsations at the 92 s WD spin period and a 3.5 h orbit (castro_arsco_2025); and CHIME/ILT J163430+445010, with a 14 min radio period and a 0.6 or 1.2 h orbital period (dong_chimefast_2025; bloot_strongly_2025).
@@
Table 2: Properties of long period transients and related Galactic sources. The sources are broadly groups according to how they are reported in the literature. We have only included published objects here, but we know of a number of new discoveries currently in preparation.
@@
Figure 6: Pulse period versus polar magnetic field ($6.4\times 10^{19}\,{\rm G}\sqrt{P\dot{P}}$) diagram for pulsars, magnetars, LPTs, and related sources, following rea_long-period_2024. We show radio pulsars as well as magnetars (stars), X-ray dim isolated neutron stars (XDINS; squares), central compact objects (CCOs, also some times referred to as ‘anti-magnetars’; pentagons) and select long-period radio pulsars (diamonds) from manchester_australia_2005. We also plot LPTs (large circles; wang_detection_2025; lee_interpulse_2025; wang_discovery_2021; caleb_emission-state-switching_2024; hurley-walker_29_2024; hurley-walker_long-period_2023; hurley-walker_radio_2022; dong_chime_2025; hurley-walker_29_2024, the radio pulsar J0311+1402 (wang_discovery_2025), and the long-period magnetar 1E 1613-5055 (esposito_swift_2011). In those cases we are assuming that the measured period is the spin period, and that the moment of inertia is that of a neutron star. We plot a range of possible ‘death lines’ for purely-dipolar (top) and highly twisted (bottom) magnetospheres, based on rea_long-period_2024, which delineate a ‘death valley’ in which radio emission is expected to cease. Finally, we show the limit where the total magnetic energy is roughly the gravitational binding energy (solid red line).

### v2

@@ 2 Radio emission mechanisms and causes of variability
where $I_{\nu}$ is the specific intensity (measured in ${\rm erg\,s}^{-1}\,{\rm cm}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-2}}\,{\rm Hz}^{-1}\,{\rm sr}^{-1}$ or equivalent), the frequency is $\nu$, $k_{B}$ is Boltzmann’s constant, and $c$ is the speed of light. For a source with flux density $S_{\nu}$ and solid angle $\Omega$, we have $I_{\nu}=S_{\nu}/\Omega$. Note that this form assumes radiation is in the Rayleigh-Jeans limit, with $h\nu\ll k_{B}T$, even if that is not actually true. For non-thermal sources the brightness temperature, $T_{B}$, is different from the actual electron temperature $T_{e}$, and typically $T_{B}\gg T_{e}$.
@@ 2.1.1 Accretion and jet ejection
AGN dominate the radio sky at flux densities above a few millijanskys (at centimetre wavelengths), and most show little variability (hovatta_statistical_2007; hodge_millijansky_2013; stewart_lofar_2016; mooley_caltech-nrao_2016; bell_murchison_2019; murphy_askap_2021), although fluctuations of a few to few tens of percent are not uncommon (e.g. falcke_radio_2001; mundell_radio_2009) and variations are both more significant and faster at higher radio frequencies (e.g. ackermann_radiogamma-ray_2011; hovatta_statistical_2007), especially for sub-classes of AGN such as blazars (max-moerbeck_ovro_2012; richards_blazars_2011; max-moerbeck_time_2014).
@@ 2.3 Stellar emission: gyrosynchrotron, plasma, and electron cyclotron maser emission
In a number of stellar systems, especially for cooler, later-type stars we can detect gyrosynchrotron emission, which is an incoherent, (usually) non-thermal process related to synchrotron emission (dulk_radio_1985; gudel_stellar_2002). Here, however, the electrons are only mildly relativistic. This leads to broad-band emission that has significant (but not 100%) circular polarisation at higher frequencies, where it is optically thin. At lower frequencies we see a typical optically-thick spectral index of $\alpha=+2.5$ and a steep spectrum on the higher frequency side, with the power-law index depending on the underlying distribution of electron energies.
@@ 3.1.1 Radio detection of gamma-ray bursts
The first GRB for which a radio afterglow was detected was GRB 970508 (frail_radio_1997). High variability was observed in the first month after the explosion, which was attributed to diffractive scintillation. Since diffractive scintillation occurs only when the source is smaller than a characteristic size, the transition from the diffractive to refractive regime could be used to measure the angular size of the fireball, determining that it was a relativistically expanding explosion. The subsequent monitoring of this event in radio (frail_450_2000) demonstrated the utility of radio observations in measuring the calorimetric properties of GRBs.
@@ 3.1.3 Short gamma-ray bursts
Neutron star mergers have long been believed to be the origin of short gamma-ray bursts (eichler_nucleosynthesis_1989; kochanek_gravitational_1993; although it can be more complicated, as in troja_nearby_2022; rastinejad_kilonova_2022). Like with long GRBs there is a collimated, relativistic jet that impacts the surrounding medium, and so many of the phenomena related to long GRBs could also take place in short GRBs (granot_gamma-ray_2014; berger_short-duration_2014).
@@ 3.1.4 Neutron star mergers
While a burst of gamma-rays was seen from GW170817/GRB170817A (goldstein_ordinary_2017), it was considerably lower luminosity than that from most GRBs. The distinction is that the observer needs to be on-axis to see a GRB, while the more general neutron star merger phenomena could be observed off-axis, as was the case with GW170817/GRB170817A. Most subsequent searches for neutron star mergers at radio wavelengths still rely on multi-messenger triggers. Given the uncertain nature of many of those triggers and the wide sky areas they point to (e.g. petrov_data-driven_2022), even triggered searches still retain some aspects of wide-field radio surveys, with many sources to consider and tight control over false positives required. With these triggered searches, non-detections can still be used to constrain the geometry, energetics, and environments of neutron star mergers (gulati_constraints_2025). But even without such a trigger, the collimated afterglow from a SGRB/neutron star merger could still be detectable by wide-field radio surveys (metzger_extragalactic_2015; hotokezaka_radio_2016; lin_detectability_2020), although as with many similar events it may be a challenge to distinguish them from other related synchrotron transients. Nonetheless, they are promising targets to understand the diversity in jet launching and how neutron star mergers are tied to their environments. If detected, followup observations including astrometry and searches for scintillation can then be used to determine the jet geometry and energetics (dobie_constraining_2020), especially with next-generation radio facilities (dobie_radio_2021; corsi_radio_2024).
@@ 3.2 Supernovae
The simple picture of steady spherical mass-loss is, in many cases, contradicted by detailed radio observations. This has been studied in detail in the case of SN 1987A in the Large Magellanic Cloud (staveley-smith_structure_1993; mccray_remnant_2016; petruk_polarized_2023), through multi-wavelength modeling of the evolving morphology and spectrum, but usually this level of detail is not possible. Instead observations can usually only constrain the flux density, and in small numbers of cases the size and rough shape (with very long baseline interferometry; e.g. demarchi_radio_2022). But even the basic lightcurve information is complex, and can show multiple radio peaks (anderson_peculiar_2017; palliyaguru_vlbi_2021; rose_late-time_2024) indicating episodic mass-loss, or broadened peaks (soderberg_radio_2005; sfaradi_dense_2024) showing aspherical ejecta. In particular, we highlight the use of unbiased surveys (stroh_luminous_2021; rose_late-time_2024) in systematically probing for unexpected emission of a large, diverse sample of SNe, and further searching for (likely) SNe that were not identified previously (gal-yam_radio_2006; dong_transient_2021).
@@
Aside from core-collapse SNe, thermonuclear SNe (Type Ia; blondin_type_2024) are of considerable interest for their elemental enrichment patterns (dwek_iron_2016) and their use in cosmology (riess_observational_1998; perlmutter_measurements_1999), but the underlying natures of their progenitors are still poorly understood (maoz_observational_2014; liu_type_2023). Radio observations can in principle distinguish between progenitor scenarios (horesh_early_2012; perez-torres_constraints_2014), constraining the presence of a non-degenerate companion. However, despite considerable investment of observational resources (e.g. panagia_search_2006; hancock_visibility_2011; chomiuk_evla_2012; margutti_inverse_2012; horesh_early_2012; perez-torres_constraints_2014) no type Ia SNe were detected for many years. This changed with the detection of SN2020eyj by kool_radio-detected_2023, whose radio properties show signs of interaction between a blast wave and the circumstellar medium (CSM), suggesting a non-degenerate companion. However, the optical/infrared signature of this SN is also somewhat unusual, showing signs of CSM interaction (Type Ia-CSM; silverman_type_2013), so it remains to be seen whether this signature is borne out by future observations of a larger sample of Ia SNe.
@@ 3.2.1 Luminous fast blue optical transients
Sitting at the intersection of core-collapse SNe and GRBs, there are expected to be a variety of relativistic explosions that would lack high-energy emission (e.g. meszaros_tev_2001). Such sources have begun to be discovered in increasing numbers by large optical surveys. The first of these from an emerging new class is AT2018cow (prentice_cow_2018; perley_fast_2019; ho_at2018cow_2019; margutti_embedded_2019; nayana_ugmrt_2021), which was notable for having bright, fast, blue optical emission along with very bright and long-lasting (sub)millimetre emission transitioning, along with quickly declining centimetre emission whose properties are not easily reconciled with the millimetre emission (margutti_embedded_2019; ho_at2018cow_2019; ho_luminous_2022) in the usual way for synchrotron explosions. These sources are called ‘(luminous) fast blue optical transients’ or (L)FBOTs, with LFBOTs those with the highest peak luminosities ($10^{44}\,{\rm erg\,s}^{-1}$) that are achieved in the shortest times (few days), although FBOTs can span a range of lower luminosities and slower timescales (e.g. drout_rapidly_2014).
@@ 3.4 Tidal disruption events
Tidal disruption events (TDEs; hills_possible_1975; rees_tidal_1988; komossa_tidal_2015; see gezari_tidal_2021 for a recent review) occur when a star is disrupted while passing within the tidal radius of a supermassive black hole (note that other TDE scenarios involving compact objects and lower-mass black holes have been proposed, e.g. krolik_swift_2011). This leads to emission potentially across the electromagnetic spectrum, including at radio wavelengths (van_velzen_radio_2011; giannios_radio_2011; see alexander_radio_2020 for a recent review).
@@ 3.5.5 Late F, G & K type main sequence stars
Finally, some F, G & K stars exhibit coherent bursts from a range of mechanisms (e.g. matthews_radio_2025) — although in general the stellar equivalents of solar radio bursts have been primarily studied on K & M dwarfs due to their extremely strong magnetic fields, since these result in stronger and more frequent flare activity (see Section 3.5.7).
@@ 3.7 X-ray binaries
Low-mass X-ray binaries, where neutron stars or black holes accrete from stellar-mass companions, are highly variable sources seen across the electromagnetic spectrum (fender_overview_2014; van_den_eijnden_new_2021; fender_jets_2006). The jets produced by these systems inject energy into the surrounding medium (gallo_dark_2005), regulating star formation and ionising the surrounding gas (fender_energization_2005), and can be used to study the formation of jets and accretion across many orders of magnitude in accretor mass (from stellar-mass systems to supermassive black holes; gallo_universal_2003; gallo_radiox-ray_2014; van_den_eijnden_new_2021), giving insight into the fundamental mechanisms of jet launching (mcclintock_black_2014) and synchrotron-emitting regions (e.g. chauhan_broadband_2021). Detailed imaging of the radio jets can also show direct evidence for relativistic jet launching and superluminal motion (mirabel_superluminal_1994; fender_ultra-relativistic_2004).
@@
Observing such systems over time and across wavelengths, there is evidence for correlation between the radio and X-ray luminosities that extends across many orders of magnitude in mass (merloni_fundamental_2003; migliari_jets_2006; gallo_universal_2003) when the X-ray emission is in the ‘hard’ state, but which may also break into multiple tracks with different levels of radio emission (soleri_nature_2011) whose origin and significance are not understood (gallo_hard_2018). However, when the X-ray emission is in other spectral states (‘intermediate’ or ‘soft’; belloni_states_2010) this correlation disappears as steady jet emission stops and discrete blobs may be ejected (fender_towards_2004). The transitions between these states likely represent disc instabilities changing the rate at which material moves from the disc onto the compact object (coriat_revisiting_2012). Such transitions can give rise to transient radio jets, either going from the ‘soft’ X-ray state to the ‘hard’ or the other way, when newly-launched jets can impact the material from previous ejection cycles (koljonen_2006_2013).
@@ 3.9 Long period transients and Galactic Centre Radio Transients
LPTs also share some radio properties with radio pulsating white dwarf binary systems (that consist of a white dwarf star with an M-dwarf companion) [^8] Although these are sometimes referred to in the literature as ‘white dwarf pulsars’, we find this terminology confusing as (a) it mixed both physical and phenomenological information, and (b) ‘pulsar’ is a well established term for neutron stars, which are not present in these systems.. There have been potentially six such systems detected to date: AR Sco, with a radio period of $\sim 1.95$ min in a 3.5 h orbit (marsh_radio-pulsing_2016); J1912-4410, with a radio period of $\sim 5.3$ min in a 4 h orbit (pelisoli_53-min-period_2023); ILT J1101++5521, with a much longer period of 2.1 h, matching the orbital period (de_ruiter_transient_2024); GLEAM-X J0704-37, with a period of 2.9 h (again matching the orbit; hurley-walker_29_2024; rodriguez_spectroscopic_2025); SDSS J230641.47+244055.8 with pulsations at the 92 s WD spin period and a 3.5 h orbit (castro_arsco_2025); and CHIME/ILT J163430+445010, with a 14 min radio period and a 0.6 or 1.2 h orbital period (dong_chimefast_2025; bloot_strongly_2025).
@@
Table 2: Properties of long period transients and related Galactic sources. The sources are broadly grouped according to how they are reported in the literature. We have only included published objects here, but we know of a number of new discoveries currently in preparation.
@@
Figure 6: Pulse period versus polar magnetic field ($6.4\times 10^{19}\,{\rm G}\sqrt{P\dot{P}}$) diagram for pulsars, magnetars, LPTs, and related sources, following rea_long-period_2024. We show radio pulsars as well as magnetars (stars), X-ray dim isolated neutron stars (XDINS; squares), central compact objects (CCOs, also some times referred to as ‘anti-magnetars’; pentagons) and select long-period radio pulsars (diamonds) from manchester_australia_2005. We also plot LPTs (large circles; wang_detection_2025; lee_interpulse_2025; wang_discovery_2021; caleb_emission-state-switching_2024; hurley-walker_29_2024; hurley-walker_long-period_2023; hurley-walker_radio_2022; dong_chime_2025; hurley-walker_29_2024), the radio pulsar J0311+1402 (wang_discovery_2025), and the long-period magnetar 1E 1613-5055 (esposito_swift_2011). In those cases we are assuming that the measured period is the spin period, and that the moment of inertia is that of a neutron star. We plot a range of possible ‘death lines’ for purely-dipolar (top) and highly twisted (bottom) magnetospheres, based on rea_long-period_2024, which delineate a ‘death valley’ in which radio emission is expected to cease. Finally, we show the limit where the total magnetic energy is roughly the gravitational binding energy (solid red line).